OpenStudy (anonymous):

Please help! I will give a medal How do I find the LCD of two rational expressions? Can you also give an example

OpenStudy (anonymous):

@phi can you help?

OpenStudy (phi):

do you have a specific problem?

OpenStudy (anonymous):

No, I just help explaining it.

OpenStudy (phi):

if you have two denominators, you would first factor them into their prime factors then you would choose the prime factors that comprise both denominators for example , if you have 1/2 and 1/4 and you want the least common denominator (which is 4) you would write 2 and 2*2 and 2*2 is the least common denominator

OpenStudy (phi):

if you had 1/6 and 1/15 6= 2*3 15= 3*5 we need a 2 and 3 (to get the 6). we need 3 (already have one) and 5 to get the 15 so the LCD is 2*3*5= 30

OpenStudy (phi):

Is that what you mean ?

OpenStudy (phi):

if it's algebra, it's the same idea , but we have more complicated expressions

OpenStudy (phi):

for example x^2+x and x^2-1 factor x^2+x = x(x+1) factor (x^2-1)= (x-1)(x+1) from x^2+x we need x(x+1) cross off the (x+1) from the 2nd denominator (because we already have it) that leaves (x-1) that we need the LCD is x(x+1)(x-1)

OpenStudy (anonymous):

Yes

OpenStudy (phi):

I am not sure how you describe this process clearly.

OpenStudy (anonymous):

I'm not sure either. My teacher does not ever put examples, he just expects us to understand what he's saying.

OpenStudy (phi):

are you doing algebra (with letters) or just numbers?

OpenStudy (phi):

Is your problem to write up an explanation on how to find the LCD or are you just trying to understand how to do it?

OpenStudy (anonymous):

I'm trying to figure out how to do it

OpenStudy (phi):

and is it with algebra or just numbers?

OpenStudy (phi):

For just numbers, say you had these two denominators (already factored) 2*3*5 3*5*5 from the first set we need 2*3*5 (that is all of them ) now cross off any of 2 , 3, or 5 from the second set \[ \cancel{3} \cdot \cancel{5} \cdot 5\] there is no 2, but we do have one 3 and one 5 we can cross off. we are left with one 5 that we need. the LCD is 2*3*5 * 5

OpenStudy (phi):

another example: 2*5*5*5*7 2*5*5*5*5*7*7 we take all the numbers from the first number: 2*5*5*5*7 now cross off any of those numbers from the second set: cross of one 2, three 5's and one 7: \[ \cancel{2}\cdot \cancel{5\cdot 5 \cdot 5} \cdot 5\cdot \cancel{7} \cdot 7\] the LCD is 2*5*5*5*7 * 5*7

OpenStudy (phi):

If you are doing algebra, we would use different examples. Is that what you want to learn?